### v1

# Community Forensics: Using Thousands of Generators to Train Fake Image Detectors

Abstract
One of the key challenges of detecting AI-generated images is spotting images that have been created by previously unseen generative models. We argue that the limited diversity of the training data is a major obstacle to addressing this problem, and we propose a new dataset that is significantly larger and more diverse than prior work. As part of creating this dataset, we systematically download thousands of text-to-image latent diffusion models and sample images from them. We also collect images from dozens of popular open source and commercial models. The resulting dataset contains 2.7M images that have been sampled from 4803 different models. These images collectively capture a wide range of scene content, generator architectures, and image processing settings. Using this dataset, we study the generalization abilities of fake image detectors. Our experiments suggest that detection performance improves as the number of models in the training set increases, even when these models have similar architectures. We also find that detection performance improves as the diversity of the models increases, and that our trained detectors generalize better than those trained on other datasets. [^1] https://jespark.net/projects/2024/community_forensics
1 Introduction
Figure 1: Performance vs. model diversity. We use images sampled from different numbers of open source latent diffusion models in the Community Forensics dataset to train fake image detectors (shown in Fig. 2(a)). As the number of models increases, so does the detector’s performance, even though these models have similar designs and same total number of images. This improvement is largest for test images from out-of-distribution generative model classes, such as pixel-based diffusion models or GANs. For each data point, we sample 10 random model subsets with 100K training images each and report the mean and standard error values.
Our ability to automatically generate realistic images is quickly outpacing our ability to detect them, potentially leading to a state of affairs in which neither humans nor machines can reliably tell real from fake. While the field of image forensics has been developing methods to address this problem, existing fake image detectors still struggle with generalization. These methods often excel at detecting images from generators that were present in their training sets, but fail when given images sampled from unseen models [105, 126, 91, 103].
A core challenge is dealing with the large amounts of variation between models. Each generator has a potentially unique combination of the architecture, loss function, and training distribution. Even seemingly minor differences in low-level image processing details, such as the ways that training images are resized or compressed, can strongly influence detection accuracy [126]. As a result of these model-specific idiosyncrasies, a generator’s images may evade detection, even when images from architecturally similar models exist in the training set. This issue has been exacerbated by the thousands of open source models that are now available online, many of which extend pretrained base models in complex ways.
(a) Systematically collected diffusion models (4763 models)
(b) Manually chosen open source models (19 models)
(c) Commercial models (11 models)
Figure 2: The Community Forensics dataset. Our dataset contains images sampled from three types of generative models. (a) We systematically download open-source latent diffusion models from a model-sharing community [124, 37]. (b) We select popular open source generators with a variety of architectures and training procedures. (c) We sample from both closed and open state-of-the-art commercial models. We present example images and their corresponding model names.
We hypothesize that the lack of diversity in training datasets is a major source of these shortcomings. Although today’s datasets often contain millions of fake images, they come from a relatively small number of generators (e.g., fewer than 20 models in any given previous work that we are aware of). As a result, this data fails to capture many sources of variation that one might encounter in the wild. These limitations also make it challenging to accurately benchmark performance, since it is easy for cues that work well on one set of generators to fail on others.
To address these problems, we propose Community Forensics, a dataset that is significantly more diverse and comprehensive than those in prior works (Fig. 2). Our dataset contains images generated by: (a) thousands of systematically downloaded open-source latent diffusion models, (b) hand-selected open source models with various architectures, and (c) state-of-the-art commercial models. We use this dataset to conduct a study of generalization in image forensics.
To acquire large numbers of models, we sample images from thousands of text-to-image diffusion models hosted on a popular model-sharing website, Hugging Face [37]. We exploit the fact that these models use a common programming library [124] and thus can be sampled in a standardized way. A large fraction of them are extensions of Stable Diffusion [102], but collectively capture a variety of common model variations, such as in the architecture, image processing, and image content. We also sample images from many other open source models, including GANs [40], autoregressive models [42], and consistency models [117, 81]. To help study how image content affects classification performance, we provide a corresponding set of real images that are designed to resemble the generated images. For example, we condition the text-to-image models using text obtained by captioning our real images.
[…]
We use this dataset to study generalization in the generated image detection problem. Our experiments support the hypothesis that increasing the diversity of generative models used in training is important for generalization. Through experiments, we find:
  • Classifiers trained on our dataset obtain strong performance, both on our newly proposed evaluations and on multiple previously-proposed benchmarks.
  • Adding more generative models improves performance. Fig. 1 demonstrates the performance of fake image detection when trained on samples from varying numbers of diffusion models. Notably, the performance improves as more models are added, even across different architectures.
  • Including diverse generative model architectures significantly improves results, since classifiers do not fully generalize between generator architectures. Likewise, the performance gain from including large numbers of images from any particular architecture is relatively marginal.
  • Standard classifiers perform well. In contrast to observations from recent work, we find that end-to-end training of classifiers based on CNNs or ViTs generalizes well, with qualitatively similar to that of other recognition problems.
2 Related work
Datasets for detecting generated images.
A number of datasets have been proposed for specifically detecting “deepfake” images containing manipulated faces [103, 105, 28, 68, 72, 64, 134]. Rather than focusing on face manipulation, we address of creating general-purpose methods that can detect images that have been directly produced by generative models. Wang et al. [126] proposed a widely-used dataset of CNN-generated images, mixing images from GANs [59, 60, 10, 132, 14, 94] with other models [104, 12, 71, 21, 11]. This work showed that forensics models generalize between generative models, providing motivation for training on large datasets of diverse generators. However, their classifier was trained on images from a single GAN and was highly sensitive to data augmentation parameters, and more recent work shows that it does not generalize to newer models [17, 91]. Ojha et al. [91] introduced a dataset of recent diffusion models and found that training a linear classifier on CLIP features [100] extracted from ProGAN-generated images performed well. Cozzolino et al. [18] extends this work by studying the performance of CLIP-based detectors on various generative models and datasets. Epstein et al. [35] simulated detecting fake images in an online way by training a detector up to a certain year and testing it on generators released after that year. Zhu et al. [133] collected 1.4M generated images from 8 different generators. These datasets, however, only consider a handful of models (less than 20 each), limiting the generalization of their detectors. We improve upon these works by collecting much more diverse generative models to improve the performance and generalization of the detector. In concurrent work, Hong et al. [50] acquires user-created images from Midjourney and CivitAI. This strategy is complementary to ours: while it aims to collect in-the-wild fake images, its distribution is centered on images that users share, and the models are not necessarily identifiable, making it challenging to rigorously analyze the dataset’s contents and to interpret experiments conducted on it.
Fingerprint-based image forensics methods.
Classic work on image forensics relied on methods based on image statistics [99] and physical constraints [57], rather than learning. A number of datasets have been created for detecting images that have been manipulated using traditional methods, such as with photo editors [29, 54, 24, 65, 88]. Recent works focus on detecting synthetic images by inspecting the generator fingerprints. Zhang et al. [131] and Marra et al. [82] proposed identifying the spatial fingerprints left by the generator to detect synthetic images. Others focus on spectral anomalies to detect synthetic images. Durall et al. [32] and Dzanic et al. [33] identified that CNN-generated images fail to reproduce certain spectral properties of real images. Corvi et al. [17] studies the frequency fingerprints of the generated images and analyzes the cross-architecture generalization of the detector. Bammey [9] uses high-frequency artifacts to detect generated images. However, these approaches may be brittle since the artifacts they rely on can be eliminated by post-processing [18]. We instead approach this problem in a data-driven manner, scaling the number of models, images, and architectures. Recent work has created ensembles of fake image classifiers [51]. In parallel, researchers have detected text generated by language models using supervised learning and heuristics [86, 56, 8, 122, 38, 106, 66, 116], which closely resemble those in visual forensics. However, no existing techniques that we are aware of aim to collect comprehensive datasets of community-created generators.
Out-of-distribution generalization.
Our work is related to the out-of-distribution recognition problem as it involves generalizing to unseen generators and image processing. A variety of approaches have been proposed for this problem, based on likelihood ratios [101, 73, 128], self-supervision [112, 87, 48, 125], internal model statistics [47, 107], temperature scaling [6, 74], and via energy-based models [31, 76, 34]. Work by Schuhmann et al. [111] and Hendrycks et al. [49] show that diverse training data and data augmentation is important to improving the robustness to out-of-distribution samples. Our results are in line with these conclusions, as we find that a diverse set of generative models and stronger augmentations improve generalization.
3 The Community Forensics Dataset
Dataset Models Images Architectures Training setup
Wang et al. [126] 11 362K GAN, Perceptual, Deepfake, … ProGAN [59] vs. LSUN [129]
Ojha et al. [91] 4 10K GAN, Perceptual, Diffusion, … ProGAN [59] vs. LSUN [129]
Epstein et al. [35] 14 570K Diffusion Diffusion vs. LAION [111]
Cozzolino et al. [18] 18 26K Diffusion LDM [102] vs. MS-COCO [75]
Synthbuster [9] 9 10K Diffusion Diffusion vs. Dresden [39]
GenImage [133] 8 1.4M Diffusion, GAN Diffusion, GAN vs. Diffusion, GAN
[…]
Table 1: Comparison with existing forensics datasets. We compare the size of the dataset with existing datasets containing identifiable generative models. We only count the number of generated images. Our dataset contains significantly more generative models than prior works. *: Only counting the unique evaluation set by Ojha et al. [91] as their dataset is based on Wang et al. [126].
To support our goal of studying generalization in generated image detection, we collect a dataset of images sampled from a wide range of models (Fig. 2). Our dataset consists of: (a) a large and systematically collected set of “in-the-wild” text-to-image latent diffusion models obtained from a model-sharing website, (b) hand-selected models from other open source architectures, and (c) closed and open state-of-the-art commercial models. We also pair these generated images with real images from other datasets. For all images in our dataset, we preserve the original image format whenever possible, without any additional compression or resizing. This is to mitigate potential bias and performance degradation in out-of-domain settings due to unwanted artifacts [43]. Our dataset contains significantly more models than previous works (Tab. 1) and spans a wider range of architectures, processing pipelines, and semantic contents.
3.1 Systematically collecting generative models
We perform our systematic collection using publicly available, open source [^2] We use the term “open source” to refer to models with public weights and source code, even if the models may be closed in some respects (e.g., private training data). models that use the Hugging Face diffusers library [124, 37] because: 1) it is a popular library for creating text-to-image models and is widely used by hobbyists, 2) thousands of such models are publicly indexed, and 3) it provides a standard interface by which we can sample images. We process them in the order of popularity, as indicated by the number of downloads. Our pipeline downloads each model and extracts relevant hyperparameters (e.g., number of diffusion steps), sampling pipeline configurations, and metadata from the model-sharing webpage [124, 37]. We sample images using a distribution of text prompts obtained from real images (Sec. 3.3). Since experiments suggest that there are diminishing returns for sampling large numbers of images from any given model, we sample a few hundred images from each one. Images with NSFW content are filtered out using a safety checker [16]. We obtain 4763 models with approximately 403 images each from this process.
While the lack of documentation in each model and the scale of data collection make it challenging to exactly characterize the model designs in this set, they are either entirely (or almost entirely) based on latent diffusion. More specifically, we categorize models as being based on latent diffusion if they perform a denoising process on a latent representation.  [^3] We note that this definition includes latent consistency models [80, 117], which are present in our dataset. Based on this criterion and the self-reported tags, all models in our systematically collected set appear to be based on latent diffusion. While pixel-based diffusion models also use the diffusers library (e.g., DeepFloyd [25]), they were incompatible with our automated generation pipeline. We record incompatible models such as these and manually sample a portion of them to construct an out-of-distribution test set (Sec. 3.4), or as manually-chosen models used for training data (Sec. 3.2).
We show examples of sampled images in Fig. 2. In Appendix D, we provide examples of models and information from their project pages. These models generate a variety of different types of images, with various types of preprocessing. For example, a large fraction of these models adapt variations of a popular pretrained latent diffusion model, Stable Diffusion [102], to different downstream applications, and use a number of adaptation strategies (e.g., using LoRA [52]). We provide the model metadata with each image to enable other possible forensics forensics applications. We discuss these in Appendix B and provide information about image and model licenses.
3.2 Collecting images from other architectures
Images from manually chosen models.
To ensure that our dataset contains a broader range of models, we manually select 19 models from public repositories and sample 40,738 images per model on average. We note that this number is itself on par with (or more than) prior datasets with identifiable generative models. We include several GANs (e.g., StyleGANs [62, 61, 63, 109], BigGAN [10], StyleSwin [130], GigaGAN [58], ProGAN [59], ProjectedGAN [108], GANsformer [53], SAN [118], and CIPS [7]), pixel-based diffusion models (e.g., GLIDE [89], ADM [27], and DeepFloyd [25]), latent diffusion models (e.g., VQ-Diffusion [44], Diffusion Transformers [96], and Latent Flow Matching [23]), and an autoregressive model (Taming Transformers [36])
Images from commercial models.
We sample 15K images from 11 commercial models using LAION-based captions to evaluate the generalization to state-of-the-art models with typically unknown architectures: DALL·E 2, 3 [92, 93], Ideogram V1, V2 [5], Midjourney V5, V6 [85], Firefly Image 2, 3 [4], FLUX.1-dev, schnell [69], and Imagen 3 [41].
3.3 Collecting real images
To help study how real images influence forensics models, we source real images from a variety of existing datasets: LAION [110], ImageNet [26], COCO [75], FFHQ [60], CelebA [77], MetFaces [61], AFHQ [15], Forchheim [45], IMD2020 [90], Landscapes HQ [115], and VISION [114]. [^4] Following common convention in visual forensics, we refer to these images as real images, even though they may be synthetic (e.g., containing graphic design). More precisely, our goal is to distinguish “AI-generated” versions of images from the originals.
3.4 Curating the evaluation set
We construct our evaluation set using the incompatible models from our automated sampling pipeline, commercial models (Sec. 3.2), and manually collected open source models. The evaluation set comprises 26K images sampled from 21 models not included in the training set. This includes our commercial models set and an additional 11K images from 10 models: Deci Diffusion V2 [121], GALIP [120], KandinskyV2.2 [113], Kvikontent [67], LCM-LoRA-SDv1.5, LCM-LoRA-SDXL, LCM-LoRA-SSD1B [81], Stable Cascade [97], DF-GAN [119], and HDiT [19], sampled using RAISE [22], ImageNet [26], FFHQ [60], and COCO [75]-based captions.
The generated images are paired with the source real data that are used to prompt the generators. However, since some of the real datasets do not have appropriate licenses for redistribution (e.g., LAION [110, 111]), we created a public version of our evaluation set by pairing the generated images with openly licensed COCO [75] and FFHQ [60] which allow redistribution for non-commercial purposes. The public version of our evaluation set will serve as an easily reproducible and shareable evaluation set that will complement our default set. We will refer to our default set as the comprehensive evaluation set. We also release the instructions to reconstruct our comprehensive set. However, note that it may not be possible to exactly reconstruct this set in the future due to link rot.
3.5 Generating images
Unconditional models are sampled until we reach the desired number of images. For class conditional models, we sample an equal number of images per class. To sample from text-conditional models, we gather prompts from multiple sources to ensure semantic diversity. We obtain captions from real images (Sec. 3.3). We either use captions that are already present in the dataset (when available), or we use BLIP [70] to generate them. The captions are then used to sample synthetic images. Some models such as GigaGAN [58] and HDiT [19] do not provide a pretrained model, so we instead use their pre-generated images. Generated images are saved in PNG format to avoid compression artifacts. However, Firefly [4] generated images are saved in JPEG format as their web UI does not allow downloading in PNG.
4 Experiments
Model Evaluation Set (mAP) Evaluation Set (Acc) Wang et al. [126] Ojha et al. [91] SB [9] GenImage [133] Ours Wang et al. [126] Ojha et al. [91] SB [9] GenImage [133] Ours Comp. Public Comp. Public Wang et al. [126] 0.897 0.696 0.516 0.642 0.535 0.600 0.714 0.527 0.508 0.533 0.509 0.517 Ojha et al. [91] 0.939 0.957 0.620 0.797 0.630 0.656 0.791 0.821 0.532 0.641 0.543 0.548 GenImage [133] 0.929 0.984 0.813 0.999 0.938 0.968 0.795 0.966 0.719 0.990 0.857 0.886 Ours 0.964 0.991 0.904 0.990 0.979 0.977 0.873 0.950 0.818 0.946 0.895 0.888 Ours - High res. 0.967 0.996 0.974 0.998 0.991 0.994 0.901 0.970 0.908 0.957 0.925 0.912
Table 2: Generalization of AI-generated image detectors across datasets. We evaluate the classifiers trained on our dataset on several datasets, including our own. We also evaluate several previously released classifiers. Our Comprehensive set (abbreviated as Comp.) pairs the generated images with original real data; the Public set pairs them with openly licensed COCO [75] and FFHQ [60] for license-compliant redistribution of the evaluation set (Sec. 3.4). We use plain CLIP-ViT-S [30, 55, 100] architecture with $224^{2}$ and $384^{2}$ (High res.) input resolutions, Wang et al. [126] and GenImage [133] use ResNet-50 [46] with $224^{2}$ input resolution, and Ojha et al. [91] uses CLIP-ViT-L with $224^{2}$ input resolution as the backbone. Our classifiers show robust performance across all evaluation sets, outperforming all baselines in out-of-distribution evaluations ([126, 91, 9] and Ours) while nearly matching GenImage [133] on its in-distribution evaluation set.
We use our dataset to conduct a study of generalization in visual forensics, asking a number of questions: (1) How well do forensics models trained on our dataset generalize to unseen models? (2) Does adding more models improve detection performance? (3) How does diversity of the training data affect performance? (4) What architectures and data augmentation schemes are most successful?
4.1 Training image forensics models
We train binary classifiers that detect generated images using our dataset to study the generalization in image forensics. We construct our training set of 5.4M images by pairing 2.7M generated images with 2.7M real images.
Training and evaluation setup.
[…]
Model architecture.
Building on prior works which mainly used CLIP-ViT [55, 30, 100] and ResNet-50 [46], we consider ViT [30] and ConvNeXt [78] pretrained models for our classifiers. We use a plain ViT-S backbone [30] pretrained on CLIP objective [55, 100] using LAION-2B [111], ImageNet 21K, and ImageNet 1K datasets [26]. We also experiment with a ConvNeXt-S model [78] pretrained on ImageNet 21K and ImageNet 1K datasets [26]. We replace the classification head with a linear layer with sigmoid activation that outputs the probability of the image being generated. Unlike prior works [18, 91] that freeze the CLIP-ViT backbone, we train the backbone end-to-end. The models are obtained through timm [13, 127] library on Hugging Face. We experiment with two input resolutions, $224^{2}$ and $384^{2}$, to evaluate the impact of the input resolution on the detector’s performance. We denote the detector with $384^{2}$ input resolution as High res. We implement the models using PyTorch [95]. The hyperparameters are detailed in Appendix C.
Data augmentation.
Prior work considered augmentations that were designed to simulate postprocessing, such as flipping, cropping, Gaussian blur, and JPEG recompression to train their detectors [91, 9, 18, 126]. We propose an augmentation scheme that extends this approach and compare it with previously proposed augmentation methods. We expand the set of augmentations to handle additional transformations that can occur in the wild, such as padding, resizing, rotation, and shear, and integrate them into a framework that can apply complex sequences of transformations. We introduce a modified version of RandAugment [20] that applies a randomly-ordered sequence of augmentations to the images. Specifically, our modified RandAugment samples a random number $n$ between 00 and $n_{\mathtt{max}}$ for each augmentation type. Then, it applies the augmentations in random order until $n$ augmentations are applied for each augmentation type to the image. We use various augmentations, including in-memory JPEG compression, random resizing with random interpolation methods, cropping, flipping, rotation, translation, shear, padding, and cutout.
Figure 3: Performance across generator types. We evaluate the classifier performance across five generator types – latent diffusion, pixel diffusion, commercial models, GANs, and other architecture (Stable Cascade [97]). Our classifiers show robust performance across all generator types, where prior works struggle to generalize.
4.2 Generalization to other datasets
We first evaluate how well classifiers trained on our dataset transfer to other benchmarks. In Table 2, we observe that our models outperform the prior works [91, 126, 133] in all evaluation sets except GenImage. This is expected since the GenImage evaluation set is a validation split of their training set; all of the generators are already seen by their classifier. On all unseen evaluation sets, our classifiers outperform all prior works. Notably, our classifiers achieve very high performance (0.991 mAP and 92.5% accuracy) on our comprehensive evaluation set with a significant margin compared to prior works. This gap in performance can be traced to our training data which incorporates a substantially richer variety of generators compared to prior works. Consequently, our classifiers demonstrate robust generalization to out-of-distribution data, where prior works often struggle.
To better illustrate the generalization of the classifiers, we show the performance per each generator type in Figure 3. We group them into five subsets: latent diffusion, pixel diffusion, commercial models, GANs, and other architecture type (Stable Cascade [97]). Our classifiers show strong performance across all generator types, unlike prior works which struggle to generalize to diverse architectures.
For the following experiments, we use our best-performing model (High res.) unless stated otherwise.
Figure 4: Performance with increasing number of images. We train a classifier with varying numbers of images from two sets: 1000 randomly chosen models and 10 popular (highly downloaded) models in the systematically collected subset. The classifier trained from 1000 random models outperforms 10 popular models in all cases. Notably, the accuracy gap is wider than that of mAP, which may suggest that having diversity in the models improves accuracy threshold calibration. We report the mean and standard error values for each data point across 4 randomly sampled subsets.
4.3 Impact of model diversity
Next, we examine the impact of the number of models in training data. We train classifiers with images sampled from 3 to $3$ generators and evaluate them (Fig. 1). To ensure that the gains are not due to simply sampling qualitatively different architectures, we only use our systematically collected latent diffusion models. We use an extended evaluation set that includes non-latent diffusion generators from our training set, which allows us to comprehensively assess the generalization capability of the classifiers trained exclusively on latent diffusion models. We find that the performance steadily increases with the number of models. However, the performance begins to flatten out beyond 1000 models, suggesting diminishing returns. Interestingly, the performance improves even on out-of-distribution architectures such as GANs and pixel-based diffusion models, even though the classifier is only trained on latent diffusion models.
(a) Impact of generator type diversity
(b) Classifier backbone comparison
Figure 5: (a) Performance and model diversity. We compare detection performance for commercial models using classifiers trained on different subsets of the dataset: the systematically collected latent diffusion models, the manually chosen models containing diverse generator types, and both. As diversity increases, so does performance. (b) Classifier backbone comparison. We compare the architectures across datasets: ours, GenImage [133], and Wang et al. [126]. Performance is similar between architectures.
In Figure 4, we vary the number of images from two sets: 1000 randomly chosen models and 10 popular models (as denoted by their number of downloads) downloaded from our systematically collected diffusion models. While the results show that the performance improves with more training images, it begins to plateau at approximately 27K images. Moreover, the classifier trained on 1000 models outperforms the 10 models in all cases, indicating that model diversity is important for strong performance. We also note that the accuracy gap is noticeably wider than that of mAP, which may suggest that model diversity is crucial in calibrating the accuracy thresholds of the classifiers.
Our experiments show that the performance improvements from increasing the number of models may plateau when they are limited to a single generator type (Fig. 1). In Figure 5(a), we show that the diversity of the generator type also plays a major role in generalization. We train classifiers on three different sets of training data: our systematically collected set, manually chosen set, and a full set consisting of both subsets. The systematic set comprises entirely of latent diffusion models, and the manual set contains numerous generator types, including GANs, latent and pixel-based diffusion, and autoregressive models (Sec. 3.2). The classifier trained on the manual set with more diverse generator types shows stronger performance compared to the one trained on the systematic set. Additionally, we find that the two sets are complementary; the performance is further improved when we train using both sets.
4.4 Analysis of design choices
We examine the impact of various design choices, including some suggested in earlier works. In particular, we investigate the choice of backbone models, freezing the backbone, semantic alignment between the real and generated data, and robustness to transformations.
Classifier backbone comparison.
We compare the performance of the classifier trained using CLIP-ViT [30, 55, 100] and ConvNeXt [78] backbones following our training procedure in Figure 5(b). We examine three datasets: ours, GenImage [133], and Wang et al. [126]. We observe similar performance between architectures across all datasets.
(a) Impact of frozen backbones
(b) Semantic alignment analysis
Figure 6: (a) Evaluating frozen backbones. Freezing the pretrained backbone, a common practice in prior works ([91, 18]), consistently decreases the performance. (b) Analyzing source and generated data alignment. We evaluate how the pairing of the real datasets affects performance. R denotes the real dataset used in training, and F indicates the source dataset used to obtain the captions for prompting the generators. The results suggest that pairing the source data (i.e., real data used to prompt the generators) with the generated images is not essential for performance.
Frozen backbone.
Prior works [91, 18] suggested using a frozen CLIP-ViT backbone for training the classifiers. We investigate this practice by training the classifiers with both frozen and unfrozen pretrained backbones, using CLIP-ViT [30, 100, 55] and ConvNeXt [78]. As shown in Figure 6(a), freezing the backbone consistently leads to poorer performance, indicating that end-to-end training is crucial to achieving high performance.
Semantic alignment.
Existing works often pair the generated images with the source dataset (i.e., the real dataset used to prompt or generate the images) arguing that misaligned data can introduce bias [126, 91, 18, 9]. We test this practice in Fig. 6(b) by examining the performance with both semantically aligned and misaligned real datasets. Specifically, we consider two real datasets: one comprised exclusively from LAION [110] and another combining ImageNet [26], MS-COCO [75], LandscapesHQ [115], Forchheim [45], VISION [114], and IMD2020 [90]. We sample our systematically collected latent diffusion models using these two sets and categorize the generated images by their source dataset. The resulting changes in the performance of all pairs were marginal, suggesting that strict alignment may not be as critical as previously believed.
Robustness to transformations.
Figure 7 illustrates the robustness of the classifiers against various transformations. Following prior works [91, 126], we test robustness to JPEG compression and Gaussian blur. Additionally, we examine robustness to rotation, resizing, padding, and shear, as they commonly occur in real-world scenarios. For padding, we randomly pad the width or height of the image with a given fraction and scale it back to the original size. Similarly, in resize, we randomly upsample or downsample the height and width of the image by a given fraction and resize it back to the original size (e.g., if fraction is $0.3$, resize the height and width to 0.7×0.7$\times$0.7 × or 1.3×1.3$\times$1.3 × and then scale it back to the original size). The results demonstrate that our models are more robust to transformations than existing models. Specifically, GenImage [133] is notably more sensitive to Gaussian blur, JPEG compression, and resizing artifacts; classifier by Ojha et al. [91] displays sensitivity to shear transforms, and the one by Wang et al. [126] performs poorly overall.
Figure 7: Robustness to various transformations. Our classifiers display robust performance across transformations. Other works generally show more sensitivity to these factors.
5 Discussion
In this paper, we studied the problem of generalizing to unseen generative models in synthetic image detection. We proposed a new dataset, Community Forensics, which contains 4803 models and 2.7M images collected from various public sources. We studied the impact of model diversity on the generalization performance of the classifiers and demonstrated that diverse data is a crucial factor in training a robust, generalizable forensics model. We trained classifiers on our dataset, demonstrated their ability to generalize on various settings, and evaluated previously proposed architecture and training practices.
While we only focus on generated image detection in our paper, our dataset may enable further forensics studies that can take advantage of our diverse array of generators. For example, in Appendix A, we show a preliminary result in identifying the generator used to synthesize a given image.
We do not intend for our dataset to be used to train classifiers that are directly used in the wild. Detecting in-the-wild synthetic images remains a challenging open problem, and detection errors can have severe consequences (e.g., falsely accusing an author of creating fake images or allowing misinformation to be certified as real). We hope that our work will serve as a stepping stone for future research in this area by providing tools and insights for studying generalization and data collection strategies.
Limitations.
While our dataset is diverse, a large portion of the data is diffusion-based, especially models fine-tuned on Stable Diffusion [102]. However, we note that Figure 1 shows that classifiers trained only on diffusion models still generalize reasonably well to other generator types as they represent different semantic content. Future work may consider collecting more diverse models, including more GANs, VQ-VAEs [123], and autoregressive models. We also note that the generative models sourced from the community may contain inappropriate content. While in many contexts it is important to detect such images (and we have removed images flagged by a NSFW detector), these models may require further scrutiny before being used in other downstream applications. Finally, although our experiments suggest that our forensics classifiers generalize to unseen models better than those of previous work, the error rates for forensics models are still too high to be used in many important applications.
Acknowledgements.
We thank the creators of the many open source models that we used to collect the Community Forensics dataset. We thank Chenhao Zheng, Cameron Johnson, Matthias Kirchner, Daniel Geng, Ziyang Chen, Ayush Shrivastava, Yiming Dou, Chao Feng, Zihao Wei, Zixuan Pan, Inbum Park, Rohit Banerjee, and Ang Cao for the valuable discussions and feedback. This research was developed with funding from the Defense Advanced Research Projects Agency (DARPA) under Contract No. HR001120C0123.
Appendix A Other applications
(a) Fake vs. real visualization
(b) Generator type visualization
Figure 8: Feature space visualization. We visualize a feature space of our trained classifier using 10% of our training data and the evaluation set. For better visibility, only a subset of our real datasets are visualized and the labels for real datasets are italicized. We observe a good separation between fake vs. real data, and between different generator types and real datasets.
Other applications, beyond the “real-or-fake” image forensics task, could potentially be supported by our dataset. In particular, a diverse array of generators and their corresponding images in our dataset may be valuable for addressing the generator attribution problem, where the goal is to identify the characteristics of the underlying generator that is responsible for synthesizing a given image.
Figure 8 presents a UMAP [83] visualization of the feature space of our trained classifier. We use the activation of the penultimate layer for visualization following Ojha et al. [91]. The feature space reveals interesting structure: GANs form a clearly separated cluster; most commercial models are distributed closely to latent diffusion models; real datasets such as LAION [110], ImageNet [26], COCO [75], and RAISE [22] are closely distributed, whereas CelebA [77], FFHQ [60], and Landscapes HQ [115] appear to be more isolated. It is important to note that these separations emerge naturally without explicit training. A targeted learning objective may further enhance these separations.
Building on the feature space observations, we use a $k$-nearest-neighbor classifier with $k\!=\!5$ using 10% of our training data to identify the generator types in our evaluation set. We separate generators as “known” (i.e., GANs, latent and pixel diffusions, and real data) and “unknown” (commercial models and Stable Cascade [97]) generator types and compute the confusion matrices as shown in Figure 9. Note that none of these generators are seen during training. Figure 9(a) demonstrates strong performance in identifying GANs, latent diffusion models, and real data. However, pixel-based diffusion models show lower performance, possibly due to their limited representation (only 3 models) in our training set. The classification result for the “unknown” set is shown in Figure 9(b). Interestingly, commercial models are predominantly classified as latent diffusion or GANs, while Stable Cascade [97] displays similarity to latent diffusion models despite their unique three-stage sampling process.
(a) Known architectures
(b) Unknown architectures
Figure 9: Generator type classification. We classify the generator type of a given image using $k$-nearest-neighbor. (a) Confusion matrix of “known” generator types. We observe high accuracy in GANs, latent diffusions, and real data. (b) Classification results on “unknown” architectures. Commercial models are predominantly classified as latent diffusion and GANs (disregarding ‘real’). Stable Cascade [97], which we categorized as Other generator type, shows similarity to latent diffusion models.
Appendix B Dataset composition
Generator licenses.
Figure 10: Histogram of model licenses in our dataset. A vast majority of the models use the CreativeML OpenRAIL-M license [1].
In Figure 10, we report the generator licenses in our dataset. Most of the models use the CreativeML OpenRAIL-M license [1].
Model Architecture License RealSource HF_pipeline_tag HF_diffusers_tag danbochman/ ccxl LatentDiff None coco,forchheim,imagenet,imd2020,laion, landscapesHQ,vision StableDiffusionXL- Pipeline StableDiffusionXL- Pipeline livingbox/ modern- style-v3 LatentDiff creativeml- openrail-m coco,forchheim,imagenet,imd2020,laion, landscapesHQ,vision StableDiffusion- Pipeline stable-diffusion DeepFloyd PixelDiff DeepFloyd-IF coco N/A N/A BigGAN GAN MIT imagenet N/A N/A
Table 3: Example model metadata. We log both the author and model names for the Hugging Face [37] models and only the model names for others. We also log the generator type (i.e., architecture), model license, source real dataset, and Hugging Face tags if available.
Latent Diff. GAN Pixel Diff. Other Models 4766 12 3 1 Percentage 99.67% 0.25% 0.06% 0.02%
Table 4: Model counts per architecture in the training set. A vast majority of the generators are latent diffusion models.
Figure 11: Number of images per generator type in the training set.
Commercial Latent Diff. GAN Pixel Diff. Other Models 11 6 2 1 1 Images 14918 6000 2000 2000 1000
Figure 12: Evaluation set composition.
Model metadata.
We show an example model metadata in Tab. 3. It contains the name of the models, their categorized architectures, licenses, source real datasets, and the Hugging Face tags if available.
Model composition.
The composition of the training set of Community Forensics is detailed in Tables 4 and 11. A vast majority of the models and generated images are latent diffusion. Figure 12 illustrates the composition of the evaluation set, which includes two variants of HDiT [19]: one trained on FFHQ [60] and another on ImageNet [26]. For computing metrics such as mAP and accuracy, these HDiT variants are treated as separate entities due to their distinct training data and model weights. However, when reporting the number of models in our dataset, we count them as a single model.
Appendix C Training settings
For training our classifiers, we use AdamW optimizer [79] with a learning rate of 2e-5, a weight decay of 1e-2, a batch size of 512, and mixed precision [84]. We use a cosine weight decay with a warmup of 20% of the total iterations. We train our models for 52K iterations using this setting. For the models in Figures 1 and 4, we employ shorter training iterations (3K) due to the computational overhead associated with training a substantial number of models for statistical analysis. We chose this number of iterations since we found that classifier performance begins to plateau with approximately this amount of training (Figure 13).
Figure 13: Impact of training iterations. The performance of the classifier plateaus beyond 3K iterations.
Appendix D Example model project page
Figure 14: Example model project page from Hugging Face [2, 37].
Figure 14 shows a project page from Hugging Face [37, 2]. We can see the tags associated with the model (e.g., Text-to-image, pipeline type, license), number of downloads, and sample images.